# GV-Rep: A Large-Scale Dataset for Genetic Variant Representation Learning

### v1

@@ 1 Introduction
Dataset for GV representation learning We assemble a large-scale GV dataset with more than 7.5 million GV records with diverse labels. This includes 155 well-labeled clinician verified GV records, serving as the anchor GVs for clinical usage (Section 5.3.1).
@@ 3.2.1 Dataset Description
ClinVar [20] ClinVar hosts genetic variants (GVs) supported by evidence and classified into four pathogenicity categories across 13,209 disease types: likely benign ($n=714,866$), benign ($n=195,030$), pathogenic ($n=143,348$), and likely pathogenic ($n=100,859$).
@@
OLIDA [24] The OLIDA database has been curated to include 1,808 high-quality bilocus variant combinations linked to 219 oligogenic diseases as a positive set [24], and 150,500 bilocus combinations from healthy individuals of The Thousand Genomes Project (1KGP) as a negative set [3].
@@ 5.1 Experiment Setup
To demonstrate the use case of our dataset, we run several experiments with four state-of-the-art pre-trained Genomic Foundation Models (GFMs): HyenaDNA(Hyena) [25], DNABERT2 [38], Nucleotide Transformer (NT), and Nucleotide Transformer v2 (NTv2) [10] [^1] For the details about the model version please refer to appendix D. During the model finetuning process, we attach a three-layer-CNN header to the frozen pretrained GFMs to aggregate information. PyTorch Lightning is used to implement the finetuning and evaluation process. NVIDIA V100 32GB are used for inference and finetuning. A total of 400 GPU hours are approximately used in total. We use Adam Optimizer [19] and set the learning rate to $1e^{-3}$.
@@ 5.2 Variant Property Prediction
We investigated the influence of context lengths on classification accuracy using the ClinVar disease classification task. Intuitively, the effect of a GV should be context-dependent. Longer contexts should facilitate better prediction of GV effects by modelling long-range interactions. For example, if a GV occurs in a gene-encoding region, it is more likely to be a pathogenic mutation based on it’s context. To test this hypothesis, we constructed a simple task to predict whether a given GV would lead to lung-related diseases [^2] For the list of lung-related diseases, see Appendix C. Figure 5 shows the AUC-ROC of fine-tuned GFMs versus sequence length. It is clear that as we decrease the context length, the prediction accuracy drops correspondingly. Moreover, there is a performance difference between the four models: while Nucleotide Transformer version 2 achieves the best performance with a context length of 1024, it is very sensitive to context length, and its accuracy drops steeply with reduced context length. In contrast, DNABERT2 and Hyena tend to be more robust to changes in context length.
@@
As shown in Table 2, the results are mixed. Overall, existing GFMs struggle with cell- and tissue- level tasks with an strong epigenetic influence: Gene-KO and sQTLs. In sQTLs, most models perform only slightly better than random guessing. In the Gene-KO task, three models converged to the same solution, suggesting that the GFMs are unlikely to provide meaningful representations, and instead, the added header learns to encode the gene-KO values. Additionally, we found that multi-species models (DNABERT2 and NT-V2) tend to perform better on the proposed tasks than models (Hyena and HT-Human) trained with only the human reference genome.
@@ 5.3 Genetic Variants Indexing
Approaches Given a set of unknown GVs $\mathcal{X}=\{x^{(n)}\}_{n=1}^{N}$, and the well annotated GVs set $\mathbb{A}=(\mathcal{X}_{A},\mathcal{Y}_{A})=\{(x^{(n)}_{A},y^{(n)}_{A})\}_{n=1}^{N_{A}}$. we could use GFMs as and encoder $\mathbb{E}_{\theta}$ and a distance function $d$ to form a Query $Q$. The query can be formed in two ways 1) An unknown variant $x_{i}$ is used as the keyword and search against in the annotated GVs set $\mathbb{A}$. They the query will be formed as $Q(\mathbb{E}_{\theta}(x_{i}),\mathbb{E}_{\theta}(\mathbb{A}),d)$, returning a list of unknown GVs which are similar to $x_{i}$. Such a clinical use case would be to better understand an unknown genetic variant that has been consistently identified in patient samples. In this scenario, clinicians may want to search to see if there are existing GVs that are similar to the unknown GV, such that it can be better categorized. 2) Set the keyword to be an annotated $x^{i}_{A}$, and search against an unknown set of GVs $\mathcal{X}$. This can then be used by genetic counsellors to prioritize unknown GVs from a patient. Here the query will be $Q(\mathbb{E}_{\theta}(x^{i}_{A}),\mathbb{E}_{\theta}(\mathcal{X}),d)$.
@@ 5.3.1 Comparison of Indexing Accuracy between Original and Finetuned GFMs
Metric We first check when querying with a GV with class label k, the optimal indexing algorithm should return GVs which are similar to the query and hereby most of the labels would be k. Here, we sample query vectors and search with the CVGV set, and for each query, we will take the top 10 results and count the number of returned GVs with the same label as the query. The percentage of the GVs with the same label as the query GV is used as the metric, indicating the query accuracy.
@@ 6 Limitations and Future Work
Despite the advances offered by the GV-Rep dataset in GV benchmarking, there are several areas ripe for future enhancement. To improve the fairness and applicability of these models, the dataset’s scope should be broadened to include sensitive attributes such as ethnicity and sex. This would facilitate an evaluation of model fairness across demographics, especially for groups typically underrepresented in existing genomic databases. Integrating epigenetic data, like DNA methylation patterns, could deepen our understanding of how these factors influence gene expression and the resulting phenotypic manifestations of genetic variants. Additionally, expanding analyses to include longer genetic contexts and cross-species genomic data could provide valuable evolutionary insights and accelerate translational research from model organisms to patient care.
@@
A robust evaluation of tokenization, pre-training data, and training regimes across a wide range of tasks will inform more effective model training strategies. Future research should also explore diverse fine-tuning approaches to enhance model accuracy and adaptability. Further development of polygenic risk scores and control vectors could lead to more personalized therapeutic strategies, dramatically increasing the dataset’s utility and impact on genomics and personalized medicine. Addressing these limitations and focusing on these future directions will be crucial for enhancing the dataset’s utility and impact on deep learning applications for GVs.
Appendix A Extended Related Work
Appendix B Additional Statistics
@@
Appendix C List of Diseases
@@
Appendix D Genomic Foundation Models used in Experiments
@@
Appendix E Ethics Statement
@@ Appendix F Licence
Appendix G Code and Data Availability

### v2

@@ 1 Introduction
Dataset for GV representation learning We assemble a large-scale GV dataset with more than 7.5 million GV records with diverse labels. This includes 155 well-labeled clinician verified GV records, serving as the anchor GVs for clinical usage (Table 3).
@@ 2 Preliminaries and Related Work
Deep Learning Model for Genomic Assay Prediction Prior to the emergence of GFMs trained on pure DNA sequences with a reconstruction objective, deep learning had already demonstrated success in predicting genome-scale sequencing assays such as CAGE, DNase-seq, and ChIP-seq. In these tasks, DNA sequences are first mapped to high-dimensional vectors, which are subsequently transformed into real-valued arrays representing the assay results. Over the years, a variety of deep learning models have been developed for these predictions. Early models, such as DeepBind [20], Basenji [22], and Basenji2 [21], utilized CNN-based architectures. More recent approaches have combined CNNs with transformers, as seen in models like Enformer [4] and Borzoi [26]. These models are not only effective in assay prediction but can also be applied to Genetic Variant Prediction. By using the predicted values or intermediate vector representations to form a latent space, scores for genetic variants can be directly extracted through dimensionality reduction techniques or a learned linear head. One challenge of training these models are that annotated dna sequence data with track information are needed for training, this potentially pose challenges to scale up the training.
@@ 3.2.1 Dataset Description
ClinVar [24] ClinVar hosts genetic variants (GVs) supported by evidence and classified into four pathogenicity categories across 13,209 disease types: likely benign ($n=714,866$), benign ($n=195,030$), pathogenic ($n=143,348$), and likely pathogenic ($n=100,859$). The genetic variants in ClinVar can be further classified by variant type as a missense variant, intronic variant, splice donor variant, or synonymous variant.
@@
OLIDA [29] The OLIDA database has been curated to include 1,808 high-quality bilocus variant combinations linked to 219 oligogenic diseases as a positive set [29], and 150,500 bilocus combinations from healthy individuals of The Thousand Genomes Project (1KGP) as a negative set [3]. Future models can leverage OLIDA to predict interactions between multiple variants and their combined effects on phenotypes.
@@ 5.1 Experiment Setup
To demonstrate the use case of our dataset, we run several experiments with four state-of-the-art pre-trained Genomic Foundation Models (GFMs): HyenaDNA(Hyena) [30], DNABERT2 [43], Nucleotide Transformer (NT), and Nucleotide Transformer v2 (NTv2) [10] [^1] For the details about the model version please refer to appendix C. During the model finetuning process, we attach a three-layer-CNN header to the frozen pretrained GFMs to aggregate information. PyTorch Lightning is used to implement the finetuning and evaluation process. NVIDIA V100 32GB are used for inference and finetuning. A total of 400 GPU hours are approximately used in total. We use Adam Optimizer [23] and set the learning rate to $1e^{-3}$.
@@ 5.2 Variant Property Prediction
We investigated the influence of context lengths on classification accuracy using the ClinVar disease classification task. Intuitively, the effect of a GV should be context-dependent. Longer contexts should facilitate better prediction of GV effects by modelling long-range interactions. For example, if a GV occurs in a gene-encoding region, it is more likely to be a pathogenic mutation based on it’s context. To test this hypothesis, we constructed a simple task to predict whether a given GV would lead to lung-related diseases [^2] For the list of lung-related diseases, see Appendix B. Figure 5 shows the AUC-ROC of fine-tuned GFMs versus sequence length. It is clear that as we decrease the context length, the prediction accuracy drops correspondingly. Moreover, there is a performance difference between the four models: while Nucleotide Transformer version 2 achieves the best performance with a context length of 1024, it is very sensitive to context length, and its accuracy drops steeply with reduced context length. In contrast, DNABERT2 and Hyena tend to be more robust to changes in context length.
@@
As shown in Table 2, the results are mixed. Overall, existing GFMs struggle with cell- and tissue- level tasks that are heavily influenced by complex regulatory mechanisms: Gene-KO and sQTLs. In sQTLs, most models perform only slightly better than random guessing. In the Gene-KO task, three models converged to the same solution, suggesting that the GFMs are unlikely to provide meaningful representations, and instead, the added header learns to encode the gene-KO values. Additionally, we found that multi-species models (DNABERT2 and NT-V2) tend to perform better on the proposed tasks than models (Hyena and HT-Human) trained with only the human reference genome.
@@ 5.3 Genetic Variants Indexing
Approaches Given a set of unknown GVs $\mathcal{X}=\{x^{(n)}\}_{n=1}^{N}$, and the well annotated GVs set $\mathbb{A}=(\mathcal{X}_{A},\mathcal{Y}_{A})=\{(x^{(n)}_{A},y^{(n)}_{A})\}_{n=1}^{N_{A}}$. we could use GFMs as an encoder $\mathbb{E}_{\theta}$ and a distance function $d$ to form a Query $Q$. The query can be formed in two ways 1) An unknown variant $x_{i}$ is used as the keyword and search against in the annotated GVs set $\mathbb{A}$. Then the query will be formed as $Q(\mathbb{E}_{\theta}(x_{i}),\mathbb{E}_{\theta}(\mathbb{A}),d)$, returning a list of known GVs which are similar to $x_{i}$. Such a clinical use case would be to better understand an unknown genetic variant that has been consistently identified in patient samples. In this scenario, clinicians may want to search to see if there are existing GVs that are similar to the unknown GV, such that it can be better categorized. 2) Set the keyword to be an annotated $x^{i}_{A}$, and search against an unknown set of GVs $\mathcal{X}$. This can then be used by genetic counsellors to prioritize unknown GVs from a patient. Here the query will be $Q(\mathbb{E}_{\theta}(x^{i}_{A}),\mathbb{E}_{\theta}(\mathcal{X}),d)$.
@@ 5.3.1 Comparison of Indexing Accuracy between Original and Finetuned GFMs
Metric We first check when querying with a GV with class label k, the optimal indexing algorithm should return GVs which are similar to the query and hereby most of the labels would be k. Here, we sample query vectors and search with the CVGV set, and for each query, we will take the top 10 results and count the number of returned GVs with the same label as the query. The percentage of the GVs with the same label as the query GV is used as the metric, indicating the Variants Indexing Accuracy.
@@ 6 Limitations and Future Work
While GV-Rep is unlikely to have a negative societal impact, caution must be exercised whenever this dataset is used to inform clinical decision-making. Applications developed using this dataset should be implemented with care to uphold the principle of human-in-the-loop for real-world applications. It is crucial that all genetic variants predicted by machine learning models are thoroughly reviewed and validated by geneticists and clinicians to ensure accurate and responsible use in informing patient care in clinical settings. The use of genetic data also presents risks to the privacy of an individual, which can be used for genetic discrimination. GV-Rep will be monitored and validated as more data is integrated, ensuring that the dataset continues to reflect the diversity of populations it aims to serve and minimizing the risk of bias over time. This vigilance will help ensure that GV-Rep remains a valuable tool for advancing personalized medicine, while safeguarding against unintended consequences.
@@
Despite the advances offered by the GV-Rep dataset in GV benchmarking, there are several areas ripe for enhancement. Future considerations include leveraging fine-mapping tools that integrate GWAS/QTL association signal strength and linkage disequilibrium (LD) information to prioritize potential causal variants. To improve the fairness and applicability of these models, the dataset’s scope should be broadened to include sensitive attributes such as ethnicity and sex. This would facilitate an evaluation of model fairness across demographics, especially for groups typically underrepresented in existing genomic databases. Integrating epigenetic data, like DNA methylation patterns, could deepen our understanding of how these factors influence gene expression and the resulting phenotypic manifestations of genetic variants. Expanding to include longer genetic contexts and cross-species pre-training could provide valuable evolutionary insights and accelerate translational research from model organisms to patient care.
@@
Conducting a thorough assessment of tokenization, pre-training data, and training regimes across a wide range of tasks will inform more effective model training strategies. Additionally, detailed evaluations of variant indexing and scaling laws across various diseases contexts will inform model generalization. Future research should explore diverse fine-tuning approaches to enhance model accuracy and adaptability. Further development of polygenic risk scores and control vectors could lead to more personalized therapeutic strategies, dramatically increasing the dataset’s utility and impact on genomics and personalized medicine. Addressing these limitations and focusing on these future directions will be crucial for enhancing the dataset’s utility and impact on deep learning applications for GVs.
Acknowledgements
The authors gratefully acknowledge the funding from research grants provided through the federal Pan-Canadian Artificial Intelligence Strategy at Vector Institute. Zehui Li is grateful to his family—Jinghai Li, Huiqing Wang, and Yiqiu Sun for their support. Valli is grateful for the constant support from all her family and friends, especially RJ.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] see Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] We don’t think there is a negative societal impact of our work
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes] See the proposed assumption and experiment in fig. 5
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] Not applicable
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See the repository for the data and code
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See the Figure 5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See the Figure 5
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section 3.2 and Section 4
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] All the sensitive information are Anonymized
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Additional Statistics
@@
Appendix B List of Diseases
@@
Appendix C Genomic Foundation Models used in Experiments
@@
Appendix D Ethics Statement
@@ Appendix E Licence
Appendix F Code and Data Availability
@@
Appendix G Dataset Documentation
We following the datasheet from [18] to document our dataset.
G.1 Motivation
For what purpose was the dataset created?
Who created the dataset (for example, which team, research group) and on behalf of which entity (for example, company, institution, organization)?
The dataset was created by the author of the paper. ZL, a student at Imperial College London and a visiting scholar at Vector Institute, and VS, a researcher at the Hospital for Sick Children and University Health Network, affiliated with the Vector Institute, were involved in its creation. The project was supervised by BW, who is affiliated with the University Health Network, the University of Toronto, and the Vector Institute.
Who funded the creation of the dataset?
Funding for research grants is provided through the federal Pan-Canadian Artificial Intelligence Strategy at Vector Institute.
G.2 Composition
What do the instances that comprise the dataset represent (for example, documents, photos, peo- ple, countries)?
Genetic Variants (GVs) Records and labels describing the effect of GVs.
How many instances are there in total (of each type, if appropriate)?
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
It is not sampled from a large set, rather it filter out GVs which does not appear on the hg38 chromosomes.
What data does each instance consist of?
each instance is a $(x,y)$ pair. Here, $x=(\text{ref},\text{alt},\text{annotation})$, and $y$ is the corresponding label indicating the class of GV or a real value quantifying the effects of the GV.
Is there a label or target associated with each instance?
Yes
Is any information missing from individual instances?
N/A
Are relationships between individual instances made explicit (for example, users’ movie ratings, social network links)?
N/A
Are there recommended data splits (for example, training, development/validation, testing)?
We use 0.8:0.1:0.1 by default, but the user can use other split with our code.
Are there any errors, sources of noise, or redundancies in the dataset?
The GVs could duplicate across different tasks, however, they will be labeled with distinct labels, so the record containing these GVs are not redundant.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (for example, websites, tweets, other datasets)?
self-contained.
Does the dataset contain data that might be considered confidential (for example, data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?
The original clinician-verified data is confidential but the anonymized version of this data is not confidential and can be distributed under the C BY-NC-SA 4.0 licence.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
No
Does the dataset identify any subpopulations (for example, by age, gender)?
No
Is it possible to identify individuals (that is, one or more natural persons), either directly or indirectly (that is, in combination with other data) from the dataset?
No
Does the dataset contain data that might be considered sensitive in any way (for example, data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
No. All the GVs are anonymized.
G.3 Collection process
What mechanisms or procedures were used to collect the data (for example, hardware appara- tuses or sensors, manual human curation, software programs, software APIs)?
Software mainly. See Section 3.2.
If the dataset is a sample from a larger set, what was the sampling strategy (for example, deterministic, probabilistic with specific sampling probabilities)?
NO
Who was involved in the data collection process (for example, students, crowdworkers, con- tractors) and how were they compensated (for example, how much were crowdworkers paid)?
N/A
Over what timeframe was the data collected?
The data was collected between Feb. 2024 to May 2024. For individual GVs, it has associated timestamp.
Were any ethical review processes conducted (for example, by an institutional review board)?
N/A
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (for example, websites)?
Mainly from third party public websites and a subset of the dataset is collected from the clinicians.
Were the individuals in question notified about the data collection?
N/A
Did the individuals in question consent to the collection and use of their data?
N/A
Has an analysis of the potential impact of the dataset and its use on data subjects (for example, a data protection impact analysis) been conducted?
N/A
G.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (for example, discretization or buck- eting, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
Yes. See Section 3.2
Was the “raw” data saved in addition to the preprocessed/cleaned/ labeled data (for example, to support unanticipated future uses)?
Yes. By default it will be saved to /root/data
Is the software that was used to preprocess/clean/label the data available?
Yes. All the preprocessing code is available.
G.5 Uses
Has the dataset been used for any tasks already?
No
Is there a repository that links to any or all papers or systems that use the dataset?
Yes. See Appendix F
What (other) tasks could the dataset be used for?
Any tasks which involving the representation learning of genetic variant will be able to take advantage of this dataset.
Is there anything about the composition of the dataset or the way it was collected and prepro- cessed/ cleaned/labeled that might impact future uses?
N/A
Are there tasks for which the dataset should not be used?
G.6 Distribution
Will the dataset be distributed to third parties outside of the entity (for example, company, institution, organization) on behalf of which the dataset was created?
Yes.
How will the dataset be distributed (for example, tarball on website, API, GitHub)?
Github
When will the dataset be distributed?
The dataset is now public accessable
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
No
G.7 Maintenance
Who will be supporting/hosting/maintaining the dataset?
The authors
How can the owner/curator/ manager of the dataset be contacted (for example, email address)?
Emails
Is there an erratum?
N/A
Will the dataset be updated (for example, to correct labeling errors, add new instances, delete instances)?
Yes, the dataset will be updated with new support for other genomic foundation models.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (for example, were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
N/A
Will older versions of the dataset continue to be supported/hosted/ maintained?
Yes, the follow-up update will mostly be the addition of other data sources.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
Yes, we recommend issuing the pull request through Github.